1. The molar solubility of magnesium carbonate in a 0.267 M ammonium carbonate solution is ______ M.

2. The molar solubility of silver phosphate in a 0.201 M silver nitrate solution is _____ M.
Answer #1

1) MgCO3    Mg2+(aq) + CO32-(aq)

let the solubility of MgCO3 is S mol/L

Thus, [Mg2+] = [CO32-] = S mol/L

Ammonium carbonate is a strong electrolyte which dissociates as,

(NH4)2CO3 2NH4+(aq) +  CO32-

Thus, [CO32-] = 0.267 M

Hence, overall [CO32-] = ( S + 0.267 ) M or mol/L

Ksp = [Mg2+] [CO32-]

= S (S + 0.267)

or, 3.5 x 10-8 = S (S + 0.267) = S2 + 0.267 S

owing to the very small Ksp vale, S2 can be neglected

or, 3.5 x 10-8 =  0.267 S

or, S = 1.31 x 10-7 mol/L

Hence, the solubility of MgCO3 is 1.31 x 10-7 mol/L